OpenStudy (anonymous):

Purvi and her friends tipped their waiter a total of $10 for dinners that cost $9, $8, $11 and $12 each. Use a percent proportion to solve for the percent tip they left the waiter. Please show work.

OpenStudy (anonymous):

The answer will be \[ \frac {10}{9+8+11+12} \times 100\% \]

OpenStudy (anonymous):

Basically \(\frac{tip}{total}=tip\%\)